Consider a portfolio consisting of the market, the risk free asset, and stock A with equal weights. Beta of this portfolio is 1.5. What is the beta of stock A?
Answer #1

Calculate the beta of the stock A as follows:

Portfolio beta = Weight *beta

Equal weight = 1/3 = 0.3333333

beta of the risk free rate = 0

Beta of the market = 1

-------------------------------------------

Stock A beta:

1.5 = (0*0.3333333) + (1*0.3333333) + (Beta of stock A * 0.3333333)

Beta of stock A = ( 1.5 - 0.3333333) / 0.3333333

Beta of stock A = 3.5